# Self-Improvement for Neural Combinatorial Optimization: Sample without Replacement, but Improvement

### v1

@@ 4.2.2 Growing nucleus
1 $\textsc{Sampled}\leftarrow\emptyset$ for $i=1,\dots,n$ do
2       $p\leftarrow(1-\frac{i-1}{n-1})\cdot p_{\text{min}}+\frac{i-1}{n-1}$ $\bm{a}_{1:T}^{(1)},\dots,\bm{a}_{1:T}^{(k)}\leftarrow$ SBS with $\pi$, beam width $k$ and nucleus $p$ $\textsc{Sampled}\leftarrow\textsc{Sampled}\cup\{\bm{a}_{1:T}^{(1)},\dots,\bm{a}_{1:T}^{(k)}\}$ Obtain $\mathbb{E}_{\pi}[f]$ estimate $\mu\leftarrow\mu(\bm{a}^{(1)},\dots,\bm{a}^{(k)})$ via (4) $\pi^{\prime}\leftarrow\pi$ for $j=1,\dots,k$ do
@@ 1 $\textsc{Sampled}\leftarrow\emptyset$ for $i=1,\dots,n$ do
                   $S\leftarrow S(\bm{a}_{1:m}^{(j)})\subseteq\{\bm{a}_{1:T}^{(1)},\dots,\bm{a}_{1:T}^{(k)}\}$  normal-$\triangleright$ sequences with same first $m$ tokens
@@ 5 Experimental Evaluation
We are pursuing two experimental goals: First, we want to evaluate the effectiveness of our method within the entire training cycle of the self-improvement strategy (see Algorithm 1), and we test it on three CO problems with different underlying network architectures: the two-dimensional Euclidean TSP, the CVRP, and the standard JSSP. Second, as the training relies on the effectiveness of the solution sampling, we want to assess the solution quality of our sampling strategy (especially with a low number of samples). We take various network checkpoints of all three problems and compare different sampling techniques to our proposed method.
@@ 5.1 Routing problems
We consider two prevalent routing problems, the two-dimensional Euclidean TSP and CVRP, where the constructive approach builds a tour by picking one node after another. For both problems, we reimplement the transformer-based architecture from the recent Bisimulation Quotienting (BQ) method (Drakulic et al., 2023). This model has nine layers that process the remaining nodes at each step. In the original work, due to its size, BQ is trained on expert trajectories from solvers with $N=100$ nodes. The authors report excellent results on the training distribution and superior generalization on larger graphs with $N\in\{200,500,1000\}$ nodes. We train the network on instances with 100 nodes in the same way and additionally with our self-improvement method, where we sample trajectories with Gumbeldore (SI GD). The setup is briefly outlined below. For implementation details and hyperparameters, please refer to Appendix D.2 and D.3. Furthermore, the ’Light Encoder Heavy Decoder’ (LEHD) by (Luo et al., 2023) follows a similar paradigm to BQ. Apart from supervised learning, they show a problem-specific way to learn on the TSP without any labeled data. We discuss this method in Appendix C.1 and demonstrate that SI GD outperforms it.
@@
Gumbeldore training
We generate to 1,000 random instances in each epoch for SI GD. Per instance, we sample 128 solutions in $n=4$ rounds with a beam width of $k=32$. For the TSP, we set the constant $\sigma$ to scale the advantages to $\sigma=0.3$, and for CVRP to $\sigma=3$. We refer to Appendix D.1 for how we obtained these values. For both problems, we start with $p_{\text{min}}=1$ (no nucleus sampling) and set $p_{\text{min}}=0.95$ after 500 epochs. Training on the generated data is performed in the same supervised way (in total $\sim$2.2k epochs for TSP and $\sim$2.7k for CVRP). Interestingly, SI GD takes fewer epochs on the TSP and more on the CVRP than SL, confirming that it is much harder to learn heuristics from scratch for the CVRP. Our method cannot access optimality information, so we compare model checkpoints by their average tour length on the validation set.
Inference and baselines
All evaluations are performed with the model trained on $N=100$. We are mainly interested in how the self-improving model performs compared to its supervised counterpart (BQ SL) and thus focus on greedy results. We also report the results using a beam search of width 16 by Drakulic et al. (2023) for coherence. We also list the results (with a beam search of width 1024) of the Attention Model (AM) (Kool et al., 2019b), which was trained with self-critical REINFORCE. We also report the results for POMO (Kwon et al., 2020) with its best inference technique (corresponding to $8N$ diversified solutions per instance of size $N$). POMO is considered the current state-of-the-art method of learning constructive heuristics from scratch. POMO performs impressively on the training distribution but has difficulties generalizing to larger instances. As our architectural setup is identical to BQ, we refer to the original work by (Drakulic et al., 2023) for comparison with various other methods.
Table 1: Results for TSP and CVRP. Gaps are computed w.r.t Concorde (Applegate et al., 2006) for TSP and HGS (Vidal, 2022) for CVRP.
@@ Results
Table 1 summarizes the optimality gaps. For the TSP, our proposed method achieves results on par with its SL counterpart on the training distribution $N=100$ and shows similarly strong generalization capabilities. For the CVRP, while the results are close to SL on $N=100$, the generalization difference becomes larger as $N$ grows, with, for example, a difference of $\sim 3\%$ on $N=1000$. We note that generalizing to larger instances means not only larger $N$ but also unseen vehicle capacities. Drakulic et al. (2023) analyze that the final performance of the model strongly depends on how an expert solution sorts the subtours. In particular, the model favors learning from solutions where the vehicle starts with subtours with the smallest remaining vehicle capacity at the end of the subtour (usually 0). Since we cannot guarantee during learning in our SI procedure that the vehicle is nearly optimally utilized in the subtours, we assume that this makes it more difficult for the model to generalize to unknown capacities. Nevertheless, the SI GD generalization results for CVRP still show the same strong dynamics as those of BQ SL, trained without any labeled data. This makes our method a promising alternative to RL-based policy gradient methods, especially due to its simplicity.
@@ 5.2 Job-Shop Scheduling Problem
Table 2: JSSP results on the Taillard benchmark set. ’GD $32\times 4$’ refers to using GD sampling with four rounds of beam width 32, using $p_{\text{min}}=0.8$. For GD $32\times 4$ and SBS 16, we report the mean across five repetitions.
@@
Model
We propose a novel pure transformer-based architecture inspired by the BQ principle Drakulic et al. (2023) to train a model with our SI GD method. We consider a JSSP instance as an unordered sequence of operations with positional information for operations belonging to the same job. We mask already scheduled operations at each construction step and process the sequence through a simple stack of pairs of attention layers. In each pair, we mask operations so that i.) in the first layer, only operations that belong to the same job can attend to each other, and ii.) in the second layer, only operations that need to run on the same machine can attend to each other. We describe the architecture in detail in Appendix D.4. To balance inference speed and model expressiveness, we settle on a latent dimension of 64, three pairs of attention layers with eight heads, and a feed-forward dimension of 256.
Datasets
All random instances are generated in the standard manner of Taillard (1993) with integer processing times in $[1,99]$. Since we train our model with SI GD, we only pre-generate a validation set of 100 instances of size $20\times 20$. We test our model on the well-known Taillard benchmark dataset (Taillard, 1993), computing optimality gaps to the best-known upper bounds. [^2] Available at http://optmizizer.com/TA.php and http://jobshop.jjvh.nl
Gumbeldore training
We train our model for 100 epochs, where for each epoch, we randomly choose a size from $\{15\times 10,15\times 15,15\times 20\}$ and sample 512 instances of the chosen size. We generate 128 solutions for each instance in 4 rounds of beam width 32 with $\sigma=0.05$. After 50 epochs, we set $p_{\text{min}}=0.95$. Similar to the routing problems, we sample 1,000 incomplete schedules (i.e., a partial sequence of jobs) of size 512 from the generated training data. We train the model to predict the next job to choose using a cross-entropy loss. Please see Appendix D.4 for more details.
@@
We test the model performance in four different ways: i.) unrolling the policy greedily, ii.) with a beam search of width 16, iii.) SBS of beam width 16 with a constant nucleus of $p=0.8$ and iv.) sampling with GD to evaluate the efficiency of GD as an inference technique. We compare our approach with L2D (Zhang et al., 2020), an RL single-agent method, and ScheduleNet (Park et al., 2022), an RL multi-agent method. Both methods represent a JSSP instance as a disjunctive graph and train Graph Neural Networks (GNN) with (variants of) PPO (Schulman et al., 2017). We also compare to the variants with 500 and 5,000 steps of L2S (Zhang et al., 2024), a strong GNN-guided improvement method that transitions between complete solutions in each step. Finally, we list the recent SPN (Corsini et al., 2024) results. This concurrent work trains a Graph Attention Network in a self-improving manner on sampled solutions (we refer to Appendix E for a methodological comparison). As noted in most previous work on neural CO, a fair comparison of inference times is challenging, as they depend on the implementation (especially using deep learning frameworks) by orders of magnitude. In particular, we report the time required to solve the instances of a given size (allowing parallelization) but compare the respective methods based on their inference type (e.g., treating greedy methods as equally computationally intensive).
@@ Results
We collect the results in Table 2. In the greedy setting, we outperform all three constructive methods, L2D, ScheduleNet, and SPN, by a wide margin and obtain smaller gaps than L2S with 500 improvement steps in all but one case. The quadratic complexity of our architecture becomes noticeable for $100\times 20$. When allowing for longer inference times, we compare our method only to the strongest method, L2S with 5,000 improvement steps. We note that our method does not always improve with beam search, with some results even worse than greedy. That sequences with high overall probability do not necessarily mean high quality is a ubiquitous effect in NLP (Holtzman et al., 2020), showing that the model is not as confident as for the routing problems. Thus, we sample with SBS and a constant nucleus of $p=0.8$. This maintains short inference times but outperforms L2S across all problem sizes. We can further improve the results by sampling with GD in multiple rounds (with $p_{\text{min}}=0.8$) and using our policy update. For more results and comparisons when using GD at inference time, see Appendix C.2.
@@ 5.3 Sampling performance
We analyze the performance of GD as a sampling technique, especially when sampling only a small number of sequences. For this, we take the model weights from two checkpoints during training: One late in training, when the model has almost converged, and an intermediate one, when the model shows a fair greedy performance but still has much room for improvement. Taking a checkpoint very early in training carries limited information, as the performance still depends strongly on weight initialization, and any sampling method with high exploration leads to substantial improvement. We then sample up to 640 sequences with an SBS beam width of 32 in up to $n=20$ rounds (for sampling with replacement, we sample the equivalent number of $32\cdot n$ sequences) and compare the optimality gap of the best sequences found. In Figure 3, we plot the results using our proposed method (’Gumbeldore’) and for sampling with (’Sample WR’) and without replacement (’Sample WOR’). To better understand the impact of the growing nucleus in Section 4.2.2, we also show the results for sampling without replacement with the growing nucleus, but without the policy update (6) (’Sample WOR (nucleus)’). We note that applying only one round of Sample WOR (nucleus) or GD is identical to Sample WOR by design. We use the same values for $p_{\text{min}}$ and $\sigma$ as during training.
In line with Shi et al. (2020), we observe that sampling WOR yields a consistent but slight improvement over WR for the routing problems but is almost unobservable for JSSP, which has a much longer sequence length (225 for $15\times 15$ and 400 for $20\times 20$). Generally, allowing different nucleus sizes can significantly improve the intermediate models while giving enough room for exploration. Further updating the policy between rounds with GD has a strong impact, especially on the intermediate models. For example, with only four rounds, we obtain an absolute improvement over Sample WOR of approx. 2% for CVRP and >1% for JSSP $20\times 20$ and reach an optimality gap with only 128 samples, which Sample WOR does not reach with 640 samples. While generally shrinking for only a few rounds, the relative improvement of GD is still significant in late models. In particular, GD consistently finds better solutions in late models when nucleus sampling only has limited influence. In Appendix C.3, we further illustrate the applicability of our method on a toy problem: we adapt our method for the board game Gomoku (Five in a Row) and evaluate how long it takes to consistently beat a deterministic expert bot.

### v2

@@ 4.2.2 Growing Nucleus
1 $\textsc{Sampled}\leftarrow\emptyset$
2 for $i=1,\dots,n$ do
3       $p\leftarrow(1-\frac{i-1}{n-1})\cdot p_{\text{min}}+\frac{i-1}{n-1}$
4       $\bm{a}_{1:T}^{(1)},\dots,\bm{a}_{1:T}^{(k)}\leftarrow$ SBS with $\pi$, beam width $k$ and nucleus $p$
5       $\textsc{Sampled}\leftarrow\textsc{Sampled}\cup\{\bm{a}_{1:T}^{(1)},\dots,\bm{a}_{1:T}^{(k)}\}$
6       Obtain $\mathbb{E}_{\pi}[f]$ estimate $\mu\leftarrow\mu(\bm{a}^{(1)},\dots,\bm{a}^{(k)})$ via (4)
7       $\pi^{\prime}\leftarrow\pi$
8       for $j=1,\dots,k$ do
@@ 7       $\pi^{\prime}\leftarrow\pi$
                   $S\leftarrow S(\bm{a}_{1:m}^{(j)})\subseteq\{\bm{a}_{1:T}^{(1)},\dots,\bm{a}_{1:T}^{(k)}\}$
                    $\triangleright$ sequences with same first $m$ tokens
@@ 5 Experimental Evaluation
We are pursuing two experimental goals: First, we want to evaluate the effectiveness of our method within the entire training cycle of the self-improvement strategy (see Algorithm 1), and we test it on three CO problems with different underlying network architectures: the two-dimensional Euclidean TSP, the CVRP, and the standard JSSP. Second, as the training relies on the effectiveness of the solution sampling, we want to assess the solution quality of our sampling strategy (especially with a low number of samples). We take various network checkpoints of all three problems and compare different sampling techniques with our proposed method.
@@ 5.1 Routing Problems
We consider two prevalent routing problems, the two-dimensional Euclidean TSP and CVRP, where the constructive approach builds a tour by picking one node after another. For both problems, we reimplement the transformer-based architecture from the recent Bisimulation Quotienting (BQ) method (Drakulic et al., 2023). This model has nine layers that process the remaining nodes at each step. In the original work, due to its size, BQ is trained on expert trajectories from solvers with $N=100$ nodes. The authors report excellent results on the training distribution and superior generalization on larger graphs with $N\in\{200,500,1000\}$ nodes. We train the network on instances with 100 nodes in the same supervised manner. In addition, we train it with two variants of our proposed self-improvement method, where we sample trajectories without replacement using (a) simple round-wise SBS (SI WOR), and (b) the proposed Gumbeldore method (SI GD). The setup is briefly outlined below. For implementation details and hyperparameters, please refer to Appendix D.2 and D.3. Furthermore, the ’Light Encoder Heavy Decoder’ (LEHD) by (Luo et al., 2023) follows a similar paradigm to BQ. Apart from supervised learning, they show a problem-specific way to learn on the TSP without any labeled data. We discuss this method in Appendix C.2 and demonstrate that SI GD outperforms it.
@@
Self-improvement training
For both SI WOR and SI GD, we generate 1,000 random instances in each epoch. Per instance, we sample 128 solutions in $n=4$ rounds with a beam width of $k=32$. For SI GD, we set the constant $\sigma$ to scale the advantages to $\sigma=0.3$ for the TSP and to $\sigma=3$ for the CVRP. We refer to Appendix D.1 for how we obtained these values. For SI GD and both routing problems, we start with $p_{\text{min}}=1$ (no nucleus sampling) and set $p_{\text{min}}=0.95$ after 500 epochs. Training on the generated data is performed in the same supervised way (in total $\sim$2.2k epochs for TSP and $\sim$2.7k for CVRP). Interestingly, SI GD takes fewer epochs on the TSP and more on the CVRP than SL, confirming that it is much harder to learn heuristics from scratch for the CVRP. Our method cannot access optimality information, so we compare model checkpoints by their average tour length on the validation set.
Table 1: Results for the routing problems, with emphasis on the BQ architecture. ’Gap’ is the optimality gap w.r.t Concorde (Applegate et al., 2006) for the TSP and HGS (Vidal, 2022) for the CVRP. ’Time’ is the time needed to solve all instances. Models are trained on instances with $N=100$ nodes and used to evaluate generalization to $N\in\{200,500,1000\}$ nodes. ’beam’ refers to beam search of given width.
@@
Inference and baselines
All evaluations are performed with the model trained on $N=100$. We are mainly interested in how the self-improving model performs compared to its supervised counterpart (BQ SL) and thus focus on greedy results. We also report the results using a beam search of width 16 by Drakulic et al. (2023) for coherence. Furthermore, we list the results of (a) Google’s OR-Tools (Perron & Furnon, 2022) as a non-learning local search algorithm; (b) the prominent Attention Model (AM) (Kool et al., 2019b) (with a beam search of width 1024), which was trained with self-critical REINFORCE and forms the basis of most constructive methods for routing; (c) its multi-decoder version MDAM (Xin et al., 2021) (with a beam search of width 50), which trains multiple diverse policies; (d) POMO (Kwon et al., 2020) with its best inference technique (corresponding to $8N$ diversified solutions per instance of size $N$). (e) SGBS (Choo et al., 2022), a pure inference technique based on beam search guided by greedy rollouts, backed by POMO with hyperparameters $(\beta,\gamma)$ which are set to $(10,10)$ for the TSP and $(4,4)$ for the CVRP. POMO is considered the current state-of-the-art method of learning constructive heuristics from scratch. POMO performs impressively on the training distribution but has difficulties generalizing to larger instances. As our architectural setup is identical to BQ, we refer to the original work by (Drakulic et al., 2023) for comparison with various other methods.
@@ Results
Table 1 summarizes the optimality gaps. Both SI WOR and SI GD show strong performance when compared to training on expert trajectories, with SI GD outperforming SI WOR. For the TSP, SI GD achieves results on par with its SL counterpart on the training distribution $N=100$ and shows similarly strong generalization capabilities. For the CVRP, while the results are close to SL on $N=100$, the generalization difference becomes larger as $N$ grows, with, for example, a difference of $\sim 3\%$ on $N=1000$. We note that generalizing to larger instances means not only larger $N$ but also unseen vehicle capacities. Drakulic et al. (2023) analyze that the final performance of the model strongly depends on how an expert solution sorts the subtours. In particular, the model favors learning from solutions where the vehicle starts with subtours with the smallest remaining vehicle capacity at the end of the subtour (usually 0). Since we cannot guarantee during learning in our SI procedure that the vehicle is nearly optimally utilized in the subtours, we assume that this makes it more difficult for the model to generalize to unknown capacities. Nevertheless, the SI GD and SI WOR generalization results for CVRP still show the same strong dynamics as those of BQ SL, trained without any labeled data. This makes our method a promising alternative to RL-based policy gradient methods, especially due to its simplicity.
@@ 5.2 Job-Shop Scheduling Problem
Model
We propose a novel pure transformer-based architecture inspired by the BQ principle (Drakulic et al., 2023) to train a model with our SI GD method. We consider a JSSP instance as an unordered sequence of operations with positional information for operations belonging to the same job. We mask already scheduled operations at each construction step and process the sequence through a simple stack of pairs of attention layers. In each pair, we mask operations so that i.) in the first layer, only operations that belong to the same job can attend to each other, and ii.) in the second layer, only operations that need to run on the same machine can attend to each other. We describe the architecture in detail in Appendix D.4. To balance inference speed and model expressiveness, we settle on a latent dimension of 64, three pairs of attention layers with eight heads, and a feed-forward dimension of 256.
Datasets
All random instances are generated in the standard manner of Taillard (1993) with integer processing times in $[1,99]$. Since we train our model only with our self-improvement method, we pre-generate a validation set of 100 instances of size $20\times 20$. We test the model on the well-known Taillard benchmark dataset (Taillard, 1993), computing optimality gaps to the best-known upper bounds. [^2] Available at http://optmizizer.com/TA.php and http://jobshop.jjvh.nl
Self-improvement training
We train the policy for 100 epochs with SI WOR and SI GD. In each epoch, we randomly choose a size from $\{15\times 10,15\times 15,15\times 20\}$ and sample 512 instances of the chosen size. We generate 128 solutions for each instance in 4 rounds of beam width 32. For SI GD, we use an advantage step size of $\sigma=0.05$ and set $p_{\text{min}}=0.95$ after 50 epochs. Similar to the routing problems, we sample 1,000 incomplete schedules (i.e., a partial sequence of jobs) of size 512 from the generated training data. We train the model to predict the next job to choose using a cross-entropy loss. Please see Appendix D.4 for more details.
[…]
@@
We test performance of the model by unrolling the policy greedily. Since SI GD shows the strongest greedy results, we further evaluate the trained model i.) with a beam search of width 16, ii.) SBS of beam width 16 with a constant nucleus of $p=0.8$ and iii.) sampling with GD to assess the efficiency of GD as an inference technique. We compare our approach with (a) L2D (Zhang et al., 2020), an RL single-agent method; (b) ScheduleNet (Park et al., 2022), an RL multi-agent method. Both methods represent a JSSP instance as a disjunctive graph and train Graph Neural Networks (GNNs) with (variants of) PPO (Schulman et al., 2017); (c) The 500- and 5,000-step variants of L2S (Zhang et al., 2024), a strong GNN-guided improvement method that transitions between complete solutions at each step; (d) the concurrent work SPN (Corsini et al., 2024), which trains a Graph Attention Network in a self-improving manner on sampled solutions (we refer to Appendix E for a methodological comparison); and (d) as a non-learning baseline, the highly efficient constraint programming solver CP-SAT in OR-Tools (Perron & Furnon, 2022), with a time limit of one hour per instance, as reported by Zhang et al. (2024). As noted in most previous work on neural CO, a fair comparison of inference times is challenging, as they depend on the implementation (especially using deep learning frameworks) by orders of magnitude. In particular, we report the time required to solve the instances of a given size (allowing parallelization) but compare the respective methods based on their inference type (e.g., treating greedy methods as equally computationally intensive).
@@ Results
We collect the results in Table 2. In the greedy setting, both SI GD and SI WOR outperform all three constructive methods, L2D, ScheduleNet, and SPN, by a wide margin and obtain smaller gaps than L2S with 500 improvement steps in all but one case. The quadratic complexity of our architecture becomes noticeable for $100\times 20$. SI GD significantly improves SI WOR. When allowing for longer inference times, we compare our strongest model SI GD only with the strongest method, L2S with 5,000 improvement steps. We note that our method does not always improve with beam search, with some results even worse than greedy. That sequences with high overall probability do not necessarily mean high quality is a ubiquitous effect in NLP (Holtzman et al., 2020), showing that the model is not as confident as for the routing problems. Thus, we use the model obtained with SI GD and sample with SBS and a constant nucleus of $p=0.8$. This maintains short inference times but outperforms L2S across all problem sizes. We can further improve the results by sampling with GD in multiple rounds (with $p_{\text{min}}=0.8$) and using our policy update. For more results and comparisons when using GD at inference time, see Appendix C.3.
@@ 5.3 Sampling Performance
We analyze the performance of GD as a sampling technique, especially when sampling only a small number of sequences. For this, we take the model weights from two checkpoints during training: One late in training, when the model has almost converged, and an intermediate one, when the model shows a fair greedy performance but still has much room for improvement. Taking a checkpoint very early in training carries limited information, as the performance still depends strongly on weight initialization, and any sampling method with high exploration leads to substantial improvement. We then sample up to 640 sequences with an SBS beam width of 32 in up to $n=20$ rounds (for sampling with replacement, we sample the equivalent number of $32\cdot n$ sequences) and compare the optimality gap of the best sequences found. In Figure 3, we plot the results using our proposed method (’Gumbeldore’) and for sampling with (’Sample WR’) and without replacement (’Sample WOR’). To better understand the impact of the growing nucleus in Section 4.2.2, we also show the results for sampling without replacement with the growing nucleus, but without the policy update (6) (’Sample WOR (nucleus)’). We note that applying only one round of Sample WOR (nucleus) or GD is identical to Sample WOR by design. We use the same value for $\sigma$ as during training, and show results for $p_{\text{min}}=1$ (no nucleus sampling) and $p_{\text{min}}=0.95$ (as during training).
In line with Shi et al. (2020), we observe that sampling WOR yields a consistent but slight improvement over WR for the routing problems but is almost unobservable for JSSP, which has a much longer sequence length (225 for $15\times 15$ and 400 for $20\times 20$). Generally, allowing different nucleus sizes can significantly improve the intermediate models while giving enough room for exploration. Updating the policy between rounds with GD has a strong impact, especially on the intermediate models. Coupling the policy update with a growing nucleus further improves the performance. For example, with only four rounds, we obtain an absolute improvement over Sample WOR of approx. 2% for CVRP and >1% for JSSP $20\times 20$ and reach an optimality gap with only 128 samples, which Sample WOR does not reach with 640 samples. While generally shrinking for only a few rounds, the relative improvement of GD is still significant in late models. In particular, GD consistently finds better solutions in late models when nucleus sampling only has limited influence. In Appendix C.4, we further illustrate the applicability of our method on a toy problem: we adapt our method for the board game Gomoku (Five in a Row) and evaluate how long it takes to consistently beat a deterministic expert bot.
@@ 7 Conclusion
Acknowledgments
We thank the Action Editor and the anonymous reviewers for their constructive feedback. This work was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - 466387255 - within the Priority Programme "SPP 2331: Machine Learning in Chemical Engineering". The authors gratefully acknowledge the Competence Center for Digital Agriculture (KoDA) at the University of Applied Sciences Weihenstephan-Triesdorf for providing computational resources.
Code Availability
Code for the experiments, data, and trained network weights are available at https://github.com/grimmlab/gumbeldore.